OpenStudy (anonymous):

Find the volume of the solid obtained by rotating the region enclosed between y=x^(1/2)+1, the line y=2 and the y-axis about the y-axis. Thanks.

OpenStudy (australopithecus):

I can take a try at this but I'm rusty so bear with me

OpenStudy (alexwee123):

cylindrical shell?

OpenStudy (australopithecus):

yeah that is what I would use

OpenStudy (alexwee123):

oh ok well it's all yours @Australopithecus :)

OpenStudy (lgbasallote):

isnt cylindrical disk useable?

OpenStudy (australopithecus):

Yeah I'm pretty sure you can do both methods

OpenStudy (australopithecus):

set 2 = x^(1/2)+1 solve for x (1)^2 = x x = 1 so we have a point of intersection at (1,2) for both functions, and the other intersection point is (0,1)  Thus, we integrate from 1 to 2 in terms of y write the function in terms of y y = x^(1/2)+1 (y+1)^2 = x then the formula for cylindrical shells is 2(pi)r(h) h = (y+1)^2 r = 2 - (y+1)^2 - 1 r = 1 - (y+1)^2 Thus it would be, integral from 1 to 2 of 2pi(1 - (y+1)^2)((y+1)^2)dy I'm probably wrong @alexwee123

OpenStudy (anonymous):

I treated it as a solid of revolution, A(y) = pi((y-1)^2)^2 = pi(y-1)^4 V = integral 2 to 1 of A(y) = pi/5 But I know if I am right?

OpenStudy (australopithecus):

@KingGeorge can you please set us straight on this question

OpenStudy (dumbcow):

@t1ger5, you are correct :) if you use disc method here you must integrate with respect to y if you use shell method , integrate with respect to x @Australopithecus you had the right idea except you put it in terms of the wrong variable